OpenStudy (harrisonshark):

Find the vertex of the parabola whose equation is y = x^2 + 8x + 12. (-4, 12) (-4, -4) (0, -6)

OpenStudy (anonymous):

clik here to find it... http://www.endmemo.com/geometry/parabola.php ans ..option B...FURTHER TAKE THE HELP FROM THIS LINK..

OpenStudy (triciaal):

vertex form of equation is f(x) = (x-h)^2 + k vertex (h, k)

OpenStudy (triciaal):

h = -b/2a = -8/2(1) = -4